OpenStudy (anonymous):

solve using row reduction: 4x+9y=1 3x+7y=1

sam (.sam.):

row reduction=elimination?

OpenStudy (anonymous):

im not exactly sure what you mean

OpenStudy (anonymous):

im trying to find x and y

OpenStudy (anonymous):

@.Sam. ,you are right

sam (.sam.):

First make x the same 4x+9y=1 times 3 3x+7y=1 times 4 ------------------------------------- 12x+27y=3 12x+28y=4 ======== -y=-1 y=1 4x+9y=1 4x+9(1)=1 4x+9=1 Solve for x

OpenStudy (anonymous):

x=1 y=-1?

jimthompson5910 (jim_thompson5910):

well row reduction deals with matrices and elimination deals with systems of equations, so they're slightly different, but they get you the same answer.

OpenStudy (anonymous):

i need to solve it using the matrix

OpenStudy (anonymous):

see here @m23chris : http://www.math.hmc.edu/calculus/tutorials/linearsystems/

OpenStudy (anonymous):

we never really did it like this its hard to understand

jimthompson5910 (jim_thompson5910):

4x+9y=1 3x+7y=1 4 9 1 3 7 1 1 9/4 1/4 (1/4)*R1 3 7 1 1 9/4 1/4 0 1/4 1/4 R2 - 3*R1 ---> R2 1 9/4 1/4 0 1 1 4*R2 1 0 -2 R1 - (9/4)*R2 ---> R1 0 1 1 So the solutions are x = -2 and y = 1

OpenStudy (anonymous):

so youre dividing by 4 in the first row

jimthompson5910 (jim_thompson5910):

yes in the third step

OpenStudy (anonymous):

ok i get it now but i have 2 more questions involving 3 rows using the matrix can you help me?

jimthompson5910 (jim_thompson5910):

sure, just make a new question and let me know about it

OpenStudy (anonymous):

ok